OpenStudy (anonymous):

How do you find the length of the transverse and conjugate axis? It would be really great if someone could explain that to me for the equation: (x-4)^2/36 - (y-2)^2/9 =1

OpenStudy (amistre64):

if you know a and b, the the length of the axis are 2a and 2b

OpenStudy (amistre64):

do you know how to draw the hyperbola of this?

OpenStudy (ranga):

Compare the given equation to the standard equation for a hyperbola: (x - h)^2 / a^2 - (y - k)^2 / b^2 = 1 You can find a and b. Then double that to get the length so f transverse and conjugate axes. Here transverse = 2a and conjugate = 2b.

OpenStudy (anonymous):

I know what the hyperbola looks like! How do you find a and b?

OpenStudy (amistre64):

the draw button is bad tonight ... the transverse axis relates to the "positive" term; underneath the positive term is a number. That number is the square(^2) of half the axis the other term is the same concept but it measures the conjugate

OpenStudy (amistre64):

(x-4)^2/36 - (y-2)^2/9 ^^^^^^^^ positive term 36 = (t/2)^2, solve for t for the transverse the other term has a 9 underneath soo... 9 = (c/2)^2, solve for c for the conjugate

OpenStudy (anonymous):

For t I got 144, is that correct?

OpenStudy (amistre64):

sqrt 36 = t/2 6*2 = t it seems like you decided to resquare the results, which is something which you do not want to do. t=12 is fine